# Robot Learning from Any Images

### v1

@@ 1 Introduction
This paper poses a fundamental question: Can we obtain robot-complete data from non-robotic images under minimal assumptions—ideally from a single image? To answer this question, we revisit the underlying reasons why prior works rely on ad hoc camera setups to build robotic environments: (1) to reconstruct the scene and object geometry as simulation assets; (2) to generate photorealistic visual observations through physics-based or image-based rendering. Our key insight is that multiview reconstruction is not necessary for building robotic scenes, as with foundation model priors, a single RGB image is sufficient for recovering a physical scene. Furthermore, high-quality visual observations can be generated without full rendering pipelines, as visual blending of real images and virtual assets can also achieve high realism. Building on these insights, we eliminate the special need for hardware setups and make real-to-sim-to-real pipelines generalizable to arbitrary single images from anywhere without requiring additional information.
@@ 2 Related Works
Physics-based Scene Generation. Several recent works [40, 41, 42, 43] explore generating a physically interactable digital twin from a single image, typically for physics-based video generation [44, 45]. However, these works generally focus on narrow object categories and show limited applications in robotics. In contrast, our approach is generalizable to diverse images and objects, and fully supports robotic simulation and data generation.
@@ 3.1 Recovering the Physical Scene from a Single Image
Our goal is to recover the entire physical scene $\{\mathcal{M},\mathcal{P},\mathbf{p},\mathbf{q},\mathcal{C}\}$ from a single image $I$. This inverse problem is inherently ill-posed. However, we demonstrate that with appropriate modeling and priors, it is possible to obtain accurate and robust estimates of the underlying scene structure.
Geometry and Appearance Modeling. We begin by obtaining textured meshes $\mathcal{M}$ for both the objects $O$ and the background $B$, which recovers the scene geometry and appearance.
For object modeling, we first generate object segmentation masks from $I$ using the grounded segmentation model [46]. The segmented object regions are then passed to a state-of-the-art image-to-3D generation model [47], which produces the object textured mesh $\mathcal{M}_{O}$. For articulated objects, since single-view articulated reconstruction remains an open problem, we retrieve similar assets from existing 3D model repositories [48, 49] as a practical alternative.
@@
We then partition $P_{S}$ into object and background point clouds, $P_{O}$ and $P_{B}$, using the object segmentation masks. A background mesh $\mathcal{M}_{B}$ is reconstructed from $P_{B}$ via surface reconstruction techniques [52]. Since the background primarily serves as a support surface in robotic tasks, we optionally approximate it using shape primitives (e.g., planes) to accelerate physics simulation.
We note that the vision foundation models [51, 46, 47, 50] exhibit strong generalization to diverse image types, enabling our method to handle in-the-wild visual data robustly.
@@
Physical Property Estimation. To enable physical interaction, we estimate key physical properties $\mathcal{P}$, such as mass (or density) and friction coefficients, required by physics simulators. To get initial estimates for different objects in a scalable way, we leverage large language models (LLMs) [54] to provide physical property estimates based on the object category and image context. Specifically, we first identify the object class, then prompt the LLM with the class name and visual context to infer plausible physical parameters. We empirically found that the estimation is quite robust to support the subsequent real-world policy deployment (see Section 4.3).
@@ 6 Limitations
This work has several limitations. (1) The quality of the generated robotic demonstrations is bounded by the fidelity of the physics simulator. However, we emphasize that the core technical contribution of this paper lies in a highly scalable and generalizable approach for transforming massive non-robotic visual data into physically plausible robotic data. While the transformed demonstrations may not match the precision of real-world collected data, our framework can generate data several orders of magnitude faster and at a much larger scale than real-world collection. By leveraging the virtually unlimited pool of Internet images, RoLA supports data generation with high diversity in object types, scene layouts, appearances, and viewpoints, significantly expanding the scalability and generalization potential of robot learning. (2) While our method does not support changing camera viewpoints within a single scene as multiview reconstruction methods do, it compensates by leveraging Internet images captured from diverse natural viewpoints. As a result, RoLA can generate training data with rich viewpoint diversity across scenes, enabling the learned policies to generalize well to novel camera perspectives.

### v2

@@ 1 Introduction
This paper poses a fundamental question: Can we obtain robot-complete data from non-robotic images under minimal assumptions—ideally from a single image? To answer this question, we revisit the underlying reasons why prior works rely on ad hoc camera setups to build robotic environments: (1) to reconstruct the scene and object geometry as simulation assets; (2) to generate photorealistic visual observations through physics-based or image-based rendering. Our key insight is that multiview reconstruction may not be necessary for building robotic scenes, as with foundation model priors, a single RGB image is sufficient for recovering a physical scene. Furthermore, high-quality visual observations can be generated without full rendering pipelines, as visual blending of real images and virtual assets can also achieve high realism. Building on these insights, we eliminate the special need for hardware setups and make real-to-sim-to-real pipelines generalizable to arbitrary single images from anywhere without requiring additional information.
@@ 2 Related Works
Physics-based Scene Generation. Several recent works [40, 41, 42, 43] explore generating a physically interactable digital twin from a single image, typically for physics-based video generation [44, 45]. However, these works generally focus on narrow object categories and show limited applications in robotics. CAST [46] is the most relevant work that also explores transforming a single image into an interactable digital twin. However, CAST doesn’t model the full background and demonstrates limited applications in robotics. In contrast, our approach is generalizable to diverse images and objects, and fully supports robotic simulation and data generation.
@@ 3.1 Recovering the Physical Scene from a Single Image
Our goal is to recover the entire physical scene $\{\mathcal{M},\mathcal{P},\mathbf{p},\mathbf{q},\mathcal{C}\}$ from a single image $I$. This inverse problem is inherently ill-posed. However, we demonstrate that with generative priors, it is possible to obtain accurate and robust estimates of the underlying scene structure.
Object and Background Modeling. We begin by obtaining textured meshes $\mathcal{M}$ for both the objects $O$ and the background $B$, which recovers the scene geometry and appearance.
For object modeling, we first generate object segmentation masks from $I$ using the grounded segmentation model [47]. The segmented object regions are then passed to a state-of-the-art image-to-3D generation model [48], which produces the object textured mesh $\mathcal{M}_{O}$.
@@
We then partition $P_{S}$ into object and background point clouds, $P_{O}$ and $P_{B}$, using the object segmentation masks. A background mesh $\mathcal{M}_{B}$ is reconstructed from $P_{B}$ via heightmap reconstruction. Since the background primarily serves as a support surface in robotic tasks, we can also approximate it using shape primitives (e.g., planes) to accelerate physics simulation.
@@
Physical Property Estimation. To enable physical interaction, we estimate key physical properties $\mathcal{P}$, such as mass (or density) and friction coefficients, required by physics simulators. To get initial estimates for different objects in a scalable way, we leverage large language models (LLMs) [52] to provide physical property estimates based on the object category and image context. Specifically, we first identify the object class, then prompt the LLM with the class name and visual context to infer plausible physical parameters.
@@ 6 Limitations
This work has several limitations. (1) The quality of the generated robotic demonstrations is bounded by the fidelity of the physics simulator. However, we emphasize that the core technical contribution of this paper lies in a highly scalable and generalizable approach for transforming massive non-robotic visual data into physically plausible robotic data. While the transformed demonstrations may not match the precision of real-world collected data, our framework can generate data several orders of magnitude faster and at a much larger scale than real-world collection. By leveraging the virtually unlimited pool of Internet images, RoLA supports data generation with high diversity in object types, scene layouts, appearances, and viewpoints, significantly expanding the scalability and generalization potential of robot learning. (2) While our method does not support changing camera viewpoints within a single scene as multiview reconstruction methods do, it compensates by leveraging Internet images captured from diverse natural viewpoints. As a result, RoLA can generate training data with rich viewpoint diversity across scenes, enabling the learned policies to generalize well to novel camera perspectives. (3) Since we rely on generative priors to recover physical scenes from partially observable scans, the generated shapes may not be well aligned with the observation and may cause collisions in physics simulation.